the width of the rectangle is one-fourth of its length. if the perimeter is 40 cm. find the dimensions of the rectangle
Answer by Mathtut(3670)

P=2L+2W or 2(L+W) so if W=1/4L and P=40 then lets substitute the knowns into our equation 40=2(L+1/4L) solving for L we get 2.5L=40 so L=16 and W=1/4(16)
resulting in W=4
answer: Length = 16
Width = 4